OpenStudy (anonymous):

Find the equation of the function. A quadratic function with x-intercepts (1,0) and (-5,0) and y-intercept (0,-5)

OpenStudy (anonymous):

@campbell_st

OpenStudy (wolf1728):

The x intercepts help you to find the 2 solutions. (x+1) (x-5) x² -4x -5 = 0 So x =-1 and x=5 The minimum value is found when the derivative is equal to zero. 2x -4=0 So the minimum is at x=2 and y= 2² -4*2 -5 y=-9

OpenStudy (wolf1728):

It seems you just need the equation and you'll just need the first 3 lines of my answer.

OpenStudy (anonymous):

I was just about to verify that! Thanks for your help.

OpenStudy (wolf1728):

u r welcome :-)